### v2

# Shapeng Jiang Lijia Wei Chen Zhang

Donald Trumps in the Virtual Polls: Predicting Public Opinions Using Large Language Modelsthanks: All authors have contributed equally and are listed in alphabetical order by last name. Corresponding author: Lijia Wei, ljwei@whu.edu.cn
Shapeng Jiang Lijia Wei Chen Zhang
In recent years, large language models (LLMs) have attracted attention due to their ability to generate human-like text. As surveys and opinion polls remain key tools for gauging public attitudes, there is increasing interest in assessing whether LLMs can accurately replicate human responses. This study examines the potential of LLMs, specifically ChatGPT-4o, to replicate human responses in large-scale surveys and to predict election outcomes based on demographic data. Employing data from the World Values Survey (WVS) and the American National Election Studies (ANES), we assess the LLM’s performance in two key tasks: predicting human survey responses and U.S. election results. In survey tasks, the LLM was tasked with generating synthetic responses for various socio-cultural and trust-related questions, demonstrating notable alignment with human response patterns across U.S.-China samples, though with some limitations on value-sensitive topics. In voting tasks, the LLM was mainly used to simulate voting behavior in past U.S. elections and predict the 2024 election outcome. Our findings show that the LLM replicates cultural differences effectively, exhibits in-sample predictive validity, and provides plausible out-of-sample forecasts, suggesting potential as a cost-effective supplement for survey-based research.
Keywords: Large language models, synthetic survey data, election prediction
Large language models (LLMs) are powerful tools for predicting human behavior due to their ability to capture the complexity of natural language and encode a wide range of human experiences, cultural norms, and decision-making patterns from extensive training data. These models reflect how people use language to express thoughts, beliefs, and emotions, making them suitable for replicating behaviors across different social contexts. This is particularly valuable in social science research, where experiments and surveys rely heavily on human participants to gather behavioral data (Kim and Lee,, 2023). Specifically, using LLMs in behavioral predictions allows researchers to scale experiments, lower the costs of human subject studies, and gain insights into individual behavior and complex social interactions, such as trust, negotiation, and cooperation (Aher et al.,, 2022). Additionally, LLMs can simulate multiple scenarios and personalities, helping researchers predict responses to new social environments (Cheng et al.,, 2023).
Previous literature has primarily focused on using large language models (LLMs) for prediction tasks, relying on training data derived from “rational” materials such as general web text, books, and articles. However, the current development trajectory of LLMs is increasingly emphasizing reasoning and rationality(Qi et al.,, 2024; Zhang et al.,, 2024), which fundamentally differs from the core requirements of public opinion prediction. Public opinion prediction inherently involves non-rational factors, such as social influences, cognitive biases, and emotional responses, which cannot be adequately captured by models trained solely on existing texts and similar materials.  [^1] As LLMs continue to evolve towards more “rational” models, the challenge remains that the presence of these non-rational elements in group decision-making will inevitably introduce errors in predictions, making it impossible to fully resolve the discrepancies between model outputs and actual public opinions.
In his book Thinking, Fast and Slow, Daniel Kahneman describes human thinking as reliant on two distinct systems: System 1, which is fast, intuitive, and influenced by tradition and experience, and System 2, which is slow, rational, and analytical (Kahneman,, 2011). This duality in human cognition is crucial for understanding the decision-making processes of individuals, particularly when they express social values or cast votes. In a similar vein, large language models (LLMs) are increasingly designed to simulate the rational, slow thinking of humans, processing vast datasets and generating responses based on reasoning and logic. However, this focus on rationality contrasts with the inherently non-rational, fast thinking that influences public opinions.
Our study introduces the Matching-LLM method [^2] The term “Matching” in the Matching-LLM method carries a dual meaning. First, it involves pairing each real human respondent with an LLM-generated counterpart. Second, it matches each real human respondent with a historical respondent from past survey data., which seeks to bridge this gap by combining the “System 2” of LLMs with the “System 1” shaped by human traditions and intuition, derived from historical data. By integrating both types of cognitive processing, we aim to more accurately model the complexities of collective human behavior, particularly in predicting public opinions. Additionally, we place a strong emphasis on incorporating historical and cultural factors into the synthesis of survey responses with LLMs. This approach acknowledges that respondents’ opinions are influenced not only by slow, rational thinking but also by the subconscious, fast thinking shaped by their regional and cultural context (Nisbett et al.,, 2001; Lodge,, 2013).
Our study is comprised of two main tasks: the survey prediction and the election prediction. For the survey task, the model adopted personas based on demographic profiles from respondents in the U.S. and China, using data from the World Values Survey (WVS). The WVS was chosen because it covers important topics related to social values and includes respondents from multiple countries, enabling cross-national comparisons. Based on these personas, the model generated synthetic responses on topics such as gender role, work ethics, trust, and morality, allowing us to compare the synthetic responses with actual human responses. For the voting task, we study the performance of LLM on predicting electoral voting. First, the model used personas from the American National Election Studies (ANES), a dataset renowned for detailed voter demographics, to recreate the outcomes of the 2016 and 2020 U.S. elections. Second, based on the voter characteristics of the 2020 U.S. presidential election, the LLM was asked to generate votes for the 2024 presidential election between Donald Trump and Kamala Harris. This prediction was completed prior to the election, and after the election results were announced, we compared the actual outcomes with the LLM’s predictions. This comparison allowed us to evaluate the model’s out-of-sample predictive capability.
To preview our results, the Matching-LLM method significantly outperforms traditional LLM approaches. The Matching-LLM effectively simulates people’s opinions on social values and captures the correlations among these questions, as well as cross-cultural differences. Additionally, the Matching-LLM voting behavior not only replicated past U.S. election outcomes but also predicted Donald Trump’s overwhelming victory in the 2024 U.S. presidential election, closely aligning with the actual results across most states. The incorporation of historical factors, which reflect human traditions and intuition derived from historical data, significantly enhanced the LLM’s performance in both tasks.
Current research on LLM-based predictions found that LLMs can accurately reproduce outcomes from classic experiments such as the ultimatum game, trust game, prisoner’s dilemma, dictator game, and strategic communication game (Aher et al.,, 2022; Xie et al.,, 2024; Phelps and Russell,, 2023; Xu et al.,, 2023; Ashokkumar et al.,, 2024). However, current LLM-based experimental predictions focus on basic experiments that are relatively simple, with clear incentives or dominant strategies. Compared to experimental predictions, surveys usually feature ambiguously defined questions with multiple options that offer minimal distinction (e.g., Likert scales). Without incentives and optimal strategies, surveys focus more on values and attitudes, where consensus may not exist.
Literature have demonstrated the ability of LLMs in predicting respondents’ choices in surveys. For instance, GPT-3, an early version of a LLM trained on data up to mid-2020, can generate responses resembling real-world survey data and capture nuanced relationships between attitudes and socio-cultural contexts (Argyle et al.,, 2023). Research also highlights the practical challenges of using LLMs in surveys. Specifically, while LLM-generated survey responses can align with real survey averages, the LLMs often struggle with variability, exhibit biases influenced by question phrasing, and homogenize responses, thereby reducing the distinctiveness observed in human data Bisbee et al., (2023); Kim and Lee, (2023); Tjuatja et al., (2023). These findings illustrate both the promise and limitations of using LLMs to simulate survey participants, underscoring the need for careful prompt design and attention to variability. In addition, these studies primarily focus on political attitudes rather than social values and are limited to predicting survey responses from English-speaking countries.
Results
WVS survey responses: in-sample prediction
Mean comparisons. Figure 1 compares human responses from the WVS sample with those generated by the LLM. Throughout this section, survey questions are referenced using concise phrases as summaries of the original questions. The figure illustrates three scenarios represented by different colors: red indicates the actual responses from human samples in WVS7, black represents the LLM-generated responses based on the demographic characteristics of the human samples in WVS7, and green depicts the weighted results, referred to as Matching-LLM, combining the LLM-generated responses (black) with historical response data from WVS6. We applied 1:1 nearest neighbor matching in the Propensity Score Matching (PSM) method to identify the most similar respondent from WVS6 for each respondent in WVS7. The responses of the matched pairs were then weighted to generate the Matching-LLM result [^3] Since Matching-LLM relies on sample information from WVS6, the results are only shown for questions common to both WVS6 and WVS7. Most of the survey questions are 3-to-10-point Likert scales..
While there are some mean differences between LLM and human responses, most LLM-synthesized or Matching-LLM-synthesized response means fall within one standard deviation of the human response means. In addition, many LLM or Matching-LLM responses exhibit a smaller variance, as their own standard deviation is largely encompassed by that of the human responses, indicating less fluctuation in LLM responses. This finding is consistent with previous research on synthetic responses(Bisbee et al.,, 2023).
Figure 1: Comparing LLM and WVS response means and standard deviations
However, in some survey items, the LLM responses diverge notably from human responses. For instance, in the U.S. sample, significant differences appear on social values questions such as “Homosexual parents”, and “Work over leisure”, which address topics like LGBTQ+ rights, and work ethic. For trust-related items (“Trust on religious outgroups” and “Trust on foreigners”), the LLM predicts a higher level of trust toward individuals from diverse religious and foreign countries than actual U.S. respondents expressed. These findings reflect that ChatGPT may exhibit a bias toward portraying Americans as more inclined toward progressive views, aligning with observations from prior studies on LLM-synthesized responses(Feng et al.,, 2023). Interestingly, for these questions, the differences between LLM and human responses are smaller in the China sample, indicating the LLM’s alignment with Chinese values and a less noticeable emphasis on socially progressive views. An exception is “Homosexual acceptance,” where the LLM still exhibits a high level of acceptance that does not align with the Chinese context.
For common knowledge items like “Question on the IMF” and “Question on the UN”, the LLM generally outperforms human respondents with minimal variation, due to its vast training data. The LLM-generated responses also tend to show less variability on moral or law questions. For example, the LLM show no tolerance for theft, bribery, and terrorism. Interestingly, when the LLM played the role of U.S. respondents, it clearly rejected parental abuse with barely any variation, yet when predicting Chinese respondents, it showed a certain level of acceptance. This result may stem from differences in parenting styles between China and the United States(Chao,, 1994), while the lack of variation in U.S. responses may be attributed to adherence to U.S. laws.
Historical Data-enhanced LLM. As shown in Figure 1, the response means of the Matching-LLM generally align more closely with the human response means in WVS7 compared to the raw LLM. However, this alignment is not necessarily guaranteed for each question, partly because the WVS is conducted every five years as repeated cross-sectional surveys, with each wave consisting of a different set of respondents, and the Matching-LLM responses are weighted using the responses of WVS6 participants who are demographically most similar to those in WVS7. Table 1 presents the mean absolute deviation (MAD) between LLM or Matching-LLM responses and human responses. Columns (1) and (4) show the differences between LLM responses and human responses for the U.S. and China samples, respectively. Similarly, columns (2) and (5) display the differences between Matching-LLM responses and human responses. The signs of these differences only indicate whether the responses generated by the LLM or Matching-LLM are greater or smaller than the human responses. Last, columns (3) and (6) present the differences between the values in columns (1) and (2), and columns (4) and (5), respectively, to examine whether the Matching-LLM approach outperforms the LLM approach.
Table 1: Matching-LLM outperforms LLM in mean comparison of responses
U.S. China
LLM Matching-LLM LLM Matching-LLM
Survey Question MAD MAD Difference MAD MAD Difference
(1) (2) (3) (4) (5) (6)
Panel A: Social values
Male-preferred employment 0.222 0.093 0.129∗∗∗ 0.056 0.010 0.046
Domestic worker priority 0.103 0.059 0.044 0.434 0.295 0.139∗∗∗
Male as breadwinner 0.240 0.148 0.092∗∗∗ 0.190 0.124 0.066∗∗
Panel B: Trust
Trust on family 0.404 0.354 0.050∗∗∗ 0.143 0.114 0.029∗∗∗
Trust on neighbors 0.195 0.166 0.029 0.014 0.015 0.001
Trust on acquaintances 0.178 0.142 0.036∗∗ 0.079 0.019 0.060∗∗∗
Trust on strangers 0.279 0.223 0.056∗∗∗ 0.046 0.012 0.034∗∗
Panel C: Ethical norms and values
Fare evasion acceptance 1.425 0.88 0.545∗∗∗ 0.200 0.404 0.204∗∗∗
Theft acceptance 0.851 0.59 0.261∗∗∗ 0.293 0.060 0.287∗∗∗
Tax evasion acceptance 0.770 0.481 0.289∗∗∗ 0.378 0.505 0.127∗∗∗
Bribery acceptance 0.717 0.42 0.297∗∗∗ 0.512 0.247 0.265∗∗∗
Homosexuality acceptance 1.482 0.873 0.609∗∗∗ 3.331 2.347 0.984∗∗∗
Divorce acceptance 0.301 0.007 0.294∗∗∗ 1.603 1.257 0.346∗∗∗
Parental abuse acceptance 0.926 0.625 0.301∗∗∗ 0.974 0.305 0.669∗∗∗
Note: *, **, and *** represent significance at 10, 5, and 1% levels, respectively.
We found that for all survey questions, the Matching-LLM responses in the U.S. sample were significantly closer to human responses compared to the LLM responses. A similar pattern was observed for the China sample overall, though with some exceptions. For instance, in the cases of “Fare evasion acceptance” and “Tax evasion acceptance,” the Matching-LLM results were significantly further from human responses. Additionally, for questions like “Male-preferred employment” and “Trust on neighbors,” the differences were not statistically significant. Given the five-year interval between the two WVS waves, we believe that the weighted results observed in China are more likely due to greater changes in social norms compared to the U.S. during this period. Overall, the results in Table 1 demonstrate that Matching-LLM significantly improves the accuracy of mean comparisons in responses compared to using LLM alone. This finding lends support to our approach of combining fast thinking (historical factors) with slow thinking (LLM).
U.S.-China differences. Figure 2 shows differences between U.S. and China responses. We observe that the LLM and Matching-LLM replicate most actual U.S.-China differences in human responses, capturing both the sign of the differences and their statistical significance. There are also some exceptions. For instance, LLM-generated U.S.-China differences in “Trust on acquaintances” and “Trust on strangers” diverge in predicted direction. Nevertheless, the Matching-LLM approach effectively corrects this divergence. In terms of common-sense knowledge, LLM responses are stable and unrelated to U.S.-China differences.
Figure 2: U.S.-China differences
Regarding questions related to ethical norms and laws, the LLM inconsistently captures U.S.-China differences, retaining them only for “Homosexuality acceptance”, “Divorce acceptance”, and “Parental abuse acceptance”. The weaker performance in these questions may stem from the LLM’s tendency to align with established laws, which can cause it to diverge from real human responses. However, the Matching-LLM preserves the U.S.-China difference in “Theft acceptance” and, for questions where the LLM simulated an incorrect direction of difference, even without correcting the direction, the Matching-LLM provides differences with smaller absolute values.
Pairwise correlation. Researchers often care about correlations between behaviors and opinions. Due to budget constraints, they often measure multiple preferences in one survey or experiment (e.g., trust and risk-taking), making the consistency of pairwise correlations among these measurements important. Therefore, the LLM should generate responses that not only have means similar to human responses but also preserve their correlations at the individual level.
We assess whether statistically significant pairwise correlations observed between two questions in the LLM-generated responses are likewise present in human responses (Snowberg and Yariv,, 2021). The results, shown in Figure 3, display the coefficients and significance of these pairwise correlations observed among questions on social values and trust. Each circle in the figure represents the correlation between a pair of questions. The horizontal axis indicates the correlation coefficients for these two questions in the human sample, while the vertical axis represents the correlation coefficients for these two questions in the LLM-generated sample (solid circles) or the Matching-LLM sample (hollow circles). Circles where pairwise correlations agree in both sign and significance between LLM-simulated and human samples are labeled as “complete agreement” and shaded red. Circles where the correlation is significant in one sample but insignificant in the other are labeled “partial disagreement” and shaded green. Lastly, circles with significant correlations in both samples but in opposite directions are marked as “complete disagreement” and shaded blue. We do not include common-sense ethical norms and values questions because LLM-simulated responses lack variation.
Figure 3: Within-subject correlations across LLM and human.
As shown in Figure 3, in the U.S. sample, only 3 out of 51 pairwise correlations indicate partial disagreement. In contrast, in the China sample, 22 out of 51 correlations show partial disagreement, while 2 exhibit complete disagreement. This reflects that the LLM-generated responses effectively preserve the correlations between various human beliefs and attitudes in the U.S. context, whereas nearly half of the correlations are not retained in China. After applying the Matching-LLM method, the correlations in the U.S. context remain consistent. Although only a subset of questions can be tested due to changes in survey items between WVS6 and WVS7, we find notable improvements within this subset in the China sample. Overall, in both the U.S. and China samples, the majority of circles fall in the first and third quadrants, indicating that the correlations between responses in the human sample are preserved by both the LLM and Matching-LLM approaches.
WVS survey responses: out-of-sample prediction
We find that using LLM alone yields reliable prediction results, and incorporating historical data for weighting further enhances the model’s ability to replicate human responses. However, thus far, the weights used in the Matching-LLM approach were optimized within this sample using a least-squares method. Specifically, the weights in the Matching-LLM approach were derived by minimizing the Euclidean distance between the Matching-LLM response vector and the human response vector. The response vector includes all the survey questions in Figure 1. Therefore, for individual questions, the Matching-LLM approach may not necessarily produce results closer to human responses than the raw LLM-generated responses. However, from the perspective of all survey questions as a whole, the Matching-LLM approach is a convex combination of LLM-generated data and human data, making its superior performance over the raw LLM approach a certainty.
To further assess the ability of the Matching-LLM approach to simulate human responses, we tested its performance on new survey questions by using the weights derived in Figure 1. In other words, we treated the survey responses simulated thus far as the training set to optimize the weights and used previously unexamined WVS survey responses as the test set to evaluate the performance of the Matching-LLM approach.
Figure 4 compares human responses with LLM-generated responses and Matching-LLM responses using the weights obtained from the training set and new WVS questions related to gender roles and family values (Method). We find that, except for the “work-motherhood impact” question in the U.S. sample, the responses generated by the Matching-LLM approach are consistently closer to human responses than those generated by the raw LLM approach. This finding reinforces confidence in the Matching-LLM approach.
Figure 4: Out-of-sample prediction: WVS survey responses.
However, these evaluations are still conducted within the WVS sample, relying on existing data to examine the predictive capabilities of the LLM and Matching-LLM approaches. In what follows, we extend our analysis to out-of-sample prediction by leveraging an upcoming real-world event—the 2024 U.S. presidential election—to test the predictive performance of both approaches.
2016 and 2020 U.S. election: in-sample prediction
We began by using data from the ANES surveys to replicate the outcomes of the 2016 and 2020 U.S. presidential elections. The LLM was instructed to adopt the persona of each ANES respondent and simulate their vote for one of the two major party candidates in the respective election. The ballots were designed to closely resemble the format used in the corresponding U.S. presidential election of that year. The individual-level results were then aggregated at the state level to determine the winner in each state, and the electoral votes were summed across states to identify the overall winner.
When comparing the LLM-generated results to the self-reported voting outcomes from ANES respondents, this constitutes an in-sample prediction. Our focus, however, is on out-of-sample prediction—comparing the LLM-generated state-level results to the actual voting outcomes. Additionally, we applied a weighting method using historical data to improve the LLM’s predictive accuracy. Specifically, we weighted the state-level aggregated results from the LLM prediction with historical voting outcomes for each state. The weights were fine-tuned until the Matching-LLM approach aligned as closely as possible with state-level outcomes from past elections.
In Figure 5, panel (a) placed the actual election results for comparison purposes, while panel (b) shows the 2016 election outcome within the ANES sample based on respondents’ self-reported voting choices, which indicates how the election would have played out using the ANES data, offering a benchmark for comparison. In panel (c), we directly prompted LLM to take the personas and casted their votes. Panel (d) displays the results generated by Matching-LLM approach, where the weight was calibrated to closely align predictions with the actual 2016 election results.
Figure 5: LLM-Simulated Replication of the 2016 U.S. Presidential Election
Due to sampling errors, the survey benchmark in panel (b) do not necessarily match the actual election outcomes. Compared with panel (a), we found that the ANES sample results differed from the actual 2016 election outcome in 8 states. Similar to ANES sample, the LLM approach incorrectly forecasted 8 states. We improved the LLM’s accuracy by incorporating historical voting data in panel (d), the Matching-LLM approach closely approximated the actual outcome, incorrectly predicting only two states: Wisconsin and Michigan.
Figure 6 reveals a pattern similar to the replication of the 2016 election. In panel (b), the 2020 ANES post-election survey underperformed, incorrectly predicting outcomes in 8 states. Similarly, the LLM prompt alone also mispredicted 8 states. However, after adjustment, the Matching-LLM results closely matched the actual outcomes, missing only two states: Arizona and North Carolina, both of which are swing states.
Figure 6: LLM-Simulated Replication of the 2020 U.S. Presidential Election
Given that Donald Trump participated in the last three U.S. presidential elections, the historical data used in the Matching-LLM approach comes from the actual results of the 2016 and 2020 elections. As a result, regardless of the chosen weight, Matching-LLM inevitably produces more accurate predictions compared to the LLM alone. Therefore, we must carefully interpret the rationale for using the Matching-LLM approach in predicting the 2016 and 2020 elections.
2024 U.S. election and 2025 German federal election: out-of-sample prediction
In this section, the purpose of employing the Matching-LLM approach is not merely to replicate known election outcomes but to derive an appropriate historical weighting scheme for predicting the 2024 U.S. presidential election and the 2025 German federal election.
To predict the 2024 election, we leveraged the demographic characteristics of respondents from ANES 2020. We derived the weights for the Matching-LLM method from in-sample election prediction, and it was then applied to forecast the outcome of the 2024 U.S. presidential election. Notably, all results in this section were obtained on or before October 30, 2024, approximately one week before Election Day on November 5, 2024. Furthermore, the corresponding working paper draft was publicly available on SSRN [^4] See https://ssrn.com/abstract=5008330. prior to the election, ensuring the predictions were made entirely ex-ante, without any retrospective adjustments based on actual results.
Using data from 6,571 respondents from the ANES 2020 survey, we predicted the outcome of the 2024 election. Both LLM and Matching-LLM methods consistently forecast that Donald Trump would win the election. Detailed state-by-state Democratic and Republican vote shares are available in Table LABEL:tab:predict_2024 in the Appendix B. Columns (5) and (6) include polling data from the website 270toWin [^5] See https://www.270towin.com/..
Figure 7: LLM-Simulated Prediction of the 2024 U.S. Presidential Election
In terms of electoral votes, both approaches produced predictions that closely matched the final outcome, with the Matching-LLM approach demonstrating superior accuracy. Donald Trump, representing the Republican Party, secured 312 electoral votes, just three votes more than the Matching-LLM prediction of 309. The standard LLM method mispredicted six states, whereas the Matching-LLM approach exhibited remarkable predictive precision, making errors in only two states: Nevada and New Hampshire. Notably, among the seven key swing states identified as crucial before the election (Arizona, Georgia, Michigan, Nevada, North Carolina, Pennsylvania, and Wisconsin), the Matching-LLM approach mispredicted only Nevada, further underscoring its effectiveness in capturing electoral dynamics.
In the forecasting of 2025 German federal election, we employed individual-level data from the CESIS database666www.gesis.org—a comprehensive repository that captures detailed demographic, socioeconomic, and political information from German voters, comprising responses from 3,965 individuals—to adjust GPT-based vote share simulations for the upcoming 2025 German federal election. By comparing the GPT-simulated results with actual historical outcomes from 2017 and 2021, we derived weighting values for each party and applied these to the 2025 predictions. After weighting the adjusted figures, our forecasts indicate that the SPD is projected to receive approximately 19.55% of the vote, while the CDU/CSU is expected to lead with about 33.11%. The Greens are forecast at 15.17%, and the FDP at 3.77%. Notably, the AfD is predicted to secure around 23.89%, and Die Linke about 4.51%. These weighted results demonstrate that our methodology effectively compensates for the systematic biases observed in the raw GPT simulations, yielding a forecast that more accurately reflects historical voting trends and provides a robust prediction for the 2025 German election.
Discussion
This study demonstrates that the latest LLM exhibit significant predictive capabilities, closely mirroring human responses when demographic variables are input. For out-of-sample predictions, the model also show potential in prediction the survey responses and US election outcomes. Additionally, based on the dual-process theory of fast and slow thinking, this study develops a method named Matching-LLM that combines historical data with synthetic data generated by the LLM to improve their predictive capabilities. Results show the Matching-LLM approach outperforms the use of LLMs alone in terms of prediction accuracy.
These findings suggest that LLMs could serve as effective supplements to traditional survey methodologies (Jansen et al.,, 2023; Rossi et al.,, 2024). By accurately predicting human responses based on demographic inputs, LLMs can reduce survey costs and serve as a supplementary data source when direct collection is limited. They also help extend survey reach, filling in data gaps for underrepresented groups. Furthermore, LLMs offer a testing ground to validate and refine survey questions, ensuring relevance to target populations. Lastly, their predictive capacity allows researchers to anticipate trends, especially in evolving social issues.
This study extends prior research on using LLMs to simulate survey responses (Bisbee et al.,, 2023; Al Tamime et al.,, 2024; Kim et al.,, 2024). First, we focused on a set of representative large-scale public opinion questions, such as work ethic, gender roles, family values, and trust. This approach is more comprehensive than similar studies and enhances confidence in using LLMs to assist large-scale survey research. Additionally, unlike prior research that focused solely on the accuracy of LLM predictions, we evaluated whether LLMs can replicate relationships between variables and across samples. We also engage with election prediction studies. Traditional election forecasts rely on macro features such as economic indicators, national opinion polls, and social media data, often analyzed using machine learning or regression models (Kennedy et al.,, 2017; Brito et al.,, 2021). For this study, limited computational resources led us to use survey data as a demographic source. However, LLMs could theoretically leverage national census data to minimize sampling errors. Moreover, ChatGPT-4o, trained on vast amounts of online data up to October 2024, offers unparalleled access to data compared to traditional methods. The primary limitation of using LLMs in election prediction, however, lies in the “black box” nature of their prediction process, which obscures the specific factors influencing voter choices.
The Matching-LLM approach proposed in this study offers a novel perspective on LLM application. While LLMs are generally recognized for effectively predicting human language and reasoning, their training on vast textual data introduces a bias toward rational thinking, limiting their ability to capture intuitive, non-rational human behavior. This motivated the development of the Matching-LLM approach, which creates a convex combination of historical data and LLM-generated data. Since human behavior is governed by both fast (intuitive) and slow (deliberative) thinking systems, this combined approach better mimics real-world human behavior than either system alone. However, this innovative approach carries risks. When mirroring human behavior, biases from slow thinking arise not only from its rational nature but also from adherence to ethical norms, laws, and progressive ideologies(Feng et al.,, 2023; Scherrer et al.,, 2024). Our results show that most deviations between LLM-simulated and human responses occur on such value-laden topics. For example, LLMs tend to be more moralistic, law-abiding, and progressive on issues like LGBTQ+ rights and gender equality. Conversely, biases in fast thinking often stem from self-reporting errors driven by social desirability bias(Phillips and Clancy,, 1972), such as income exaggeration. For instance, in the 2024 U.S. presidential election, discrepancies between opinion polls and the final results—where many polls predicted Harris’s victory—might have been due to sampling biases or reluctance among Trump supporters to participate in surveys (Boyle et al.,, 2023; Graefe,, 2024). When both thinking systems produce errors, the direction of these errors determines the efficacy of the convex combination. Opposing biases may cancel each other out, while aligned biases could amplify inaccuracies.
While this study demonstrates the potential of LLMs in predicting human responses, areas for further exploration remain. Our analysis primarily used ChatGPT-4o, without examining other LLMs that may exhibit different biases on value-laden topics due to varied training data. Moreover, our data relied on WVS and ANES surveys, and incorporating more diverse sources could improve generalizability. Lastly, our cross-cultural analysis was centered on the U.S. and China. Future studies should expand this scope to evaluate LLM performance across a broader range of cultural contexts.
Methods
We confirm that our research adheres to all applicable ethical guidelines. As the datasets used were derived from publicly available sources, participant compensation does not apply to this study. Additionally, none of the studies involved pre-registration.
Datasets
The LLM model, ChatGPT-4o, was trained on data up to October 2023, and we relied exclusively on the ChatGPT API tool, without incorporating any real-time search capabilities. In the survey task, when using LLM approach, we provided ChatGPT with demographic characteristics of respondents from the United States ($N=2077$) and China ($N=2001$), as recorded in the 7th Wave of the World Values Survey (WVS). After adopting a persona defined by a set of demographic characteristics of a respondent, the LLM was tasked with generating a dataset of synthetic opinions on social values, trust, common-sense questions, and ethical norms and values.
In the election task, when using LLM approach, the LLM was provided with the demographic characteristics of respondents from ANES2020 ($N=6571$). We selected ANES2020 for three main reasons: first, its sample siz is about twice that of ANES2016, making it more representative; second, it includes information on respondents’ voting choices in the 2016 election; and third, as of the 2024 U.S. presidential election, ANES2024 had only released a pilot study with approximately 1,900 respondents, which was insufficient for analysis.
The LLM simulated voting behavior by adopting the persona of an ANES respondent and casting votes in three elections: Hillary Clinton vs. Donald Trump (2016), Joe Biden vs. Donald Trump (2020), and Kamala Harris vs. Donald Trump (2024). Following U.S. election rules, the simulated votes were aggregated by state using the survey’s sampling weights to ensure representativeness. These state-level results were then combined nationally to calculate electoral votes and determine the winner. This process allows us to compare the LLM-generated synthetic voting results with the actual outcomes of the elections.
Matching and weights.
When employing the Matching-LLM approach, we first performed matching and then determined a unified weight for combining historical data with LLM-generated data. The methods for matching and weighting differ between predicting WVS survey responses and U.S. election voting behavior.
The Matching-LLM method consists of three steps:
  1. 1.
    Matching: This step involves using historically collected individual characteristic data to match with the features of current respondents. The goal is to generate a corresponding “System 1” response for each respondent. If the data structure follows a panel format, the historical response data is directly matched.
  2. 2.
    LLM: In this step, a large language model is employed to generate the “System 2” response for each respondent. All simulated responses generated by the LLM were obtained via the API rather than the standard visual interface.
  3. 3.
    Weighting: The final step involves weighting the responses from System 1 and System 2. For in-sample predictions, the weights are optimized to achieve the best possible results. For out-of-sample predictions, the weighting is determined by referencing the weights used for similar question types in the in-sample setting. The magnitude and application of these weightings depend on the data structure and the specific research question. For instance, when working with panel data, historical data are typically assigned a higher weight because they provide a richer and more consistent set of relevant information. In contrast, for non-panel data, the weighting applied to historical data tends to be significantly lower. Additionally, it is essential to consider the potential for polarization in ChatGPT’s responses on certain topics.
As a example, for the prediction of WVS survey responses, we begin with LLM-synthesized responses generated based on WVS7 individual characteristics. Using propensity score matching (PSM), we identify the most similar individual from WVS6 for each individual in WVS7 through 1:1 nearest-neighbor matching. The demographic variables used for matching include age, gender, education level, marital status, occupation, and income scale.
We then combine the LLM-synthesized responses from WVS7 and the responses of matched individuals from WVS6 using the following equation:
$R_{k}^{\text{Matching-LLM}}=hR_{k}^{\text{WVS6}}+(1-h)R_{k}^{\text{WVS7-LLM}}$
where $R_{k}^{\text{Matching-LLM}}$ is the response vector generated using the Matching-LLM approach for the $k$-th question, $R_{k}^{\text{WVS6}}$ represents the response vector from individuals matched in WVS6, and $R^{\text{WVS7-LLM}}$ contains the LLM-synthesized responses from WVS7. $h\in[0,1]$ denotes the weight.
We employ the least squared method to estimate the weight $\hat{h}$:
$\hat{h}=argmin\frac{1}{K}\sum_{k}\sqrt{(R_{k}^{\text{Matching-LLM}}-R_{k}^{\text{WVS7}})^{T}\cdot(R_{k}^{\text{Matching-LLM}}-R_{k}^{\text{WVS7}})}$
such that the average Euclidean distance between the response vectors $R_{k}^{\text{Matching-LLM}}$ and $R_{k}^{\text{WVS7}}$ is minimized, where $R_{k}^{\text{WVS7}}$ denotes the actual responses from WVS7. We conducted the matching and weight generation processes separately for the China and U.S. samples, resulting in weights $\hat{h}=0.31$ for the China sample and $\hat{h}=0.23$ for the U.S. sample.
For the prediction of U.S. election using the ANES2020 sample, since our focus is on state-level prediction accuracy, we simply weight the actual vote shares of the two-party candidates at the state level (from the 2016 and 2020 elections) and the LLM-synthesized vote shares. Specifically, the predicted vote share using Matching-LLM approach for party $p$ in state $i$ for year $y$ is defined as follows:
$\text{Predicted vote share}_{i,y,p}=h\times\text{Historical vote share}_{i,p}+(1-h)\times\text{LLM-predicted vote share}_{i,y,p}$
where $h$ denotes the weight. We fine-tuned $h$ until the predicted outcomes aligned as closely as possible with the actual state-level winners. We then applied this optimized weighting parameter $\hat{h}$ to forecast the outcome of the 2024 presidential election.
GPT-4o prompt
The prompt used in the LLM approach is shown below, with the demographics used detailed in the appendix.
LLM prompt: WVS
Step 1: Assigning a persona to the LLM
“It is the year 2017. You are a [AGE]-year-old [GENDER] American/Chinese living in [STATE/PROVINCE], with [EDUCATION LEVEL]. Your marital status is [MARITAL STATUS], and you [OCCUPATION DESCRIPTION]. On an income scale on which 1 indicates the lowest income group and 10 the highest income group in your country, your household is [INCOME LEVEL].”
Step 2: Setting the scenario
“Hello. I am from the World Values Survey Association. We are carrying out a global study of what people value in life. This study will interview samples representing most of the world’s people. Your name has been selected at random as part of a representative sample of the people in America. I’d like to ask your views on a number of different subjects. Your input will be treated strictly confidential, but it will contribute to a better understanding of what people all over the world believe and want out of life.”
Step 3: Asking questions and gathering responses
  1. 1.
    Social values
    “How would you feel about the following statements? Do you agree or disagree with them? Choose 1 for Agree strongly, 2 for Agree, 3 for Neither agree nor disagree, 4 for Disagree, and 5 for Disagree strongly. ”
    • Male-preferred employment. When jobs are scarce, men should have more right to a job than women
    • Domestic worker priority. When jobs are scarce, employers should give priority to people of this country over immigrants
    • Male as breadwinner. If a woman earns more money than her husband, it’s almost certain to cause problems
    • Homosexual parents. Homosexual couples are as good parents as other couples
    • Parenthood as social duty. It is a duty towards society to have children
    • Filial duty. Adult children have the duty to provide long-term care for their parents
    • Anti-laziness. People who don’t work turn lazy
    • Work as social duty. Work is a duty towards society
    • Work over leisure. Work should always come first, even if it means less spare time
  2. 2.
    Trust
    “I’d like to ask you how much you trust people from various groups. Could you tell me for each whether you trust people from this group completely, somewhat, not very much or not at all? Choose 1 for Trust completely, 2 for Trust somewhat, 3 for Do not trust very much, 4 for Do not trust at all. ”
    • Trust on family. Your family
    • Trust on neighbors. Your neighborhood
    • Trust on acquaintances. People you know personally
    • Trust on strangers. People you meet for the first time
    • Trust on religious outgroups. People of another religion
    • Trust on foreigners. People of another nationality
  3. 3.
    Common-sense on international organization
    “Here are some questions about international organizations. Many people don’t know the answers to these questions, but if you do please tell me.
    • Question of the UN. Five countries have permanent seats on the Security Council of the United Nations. Which one of the following is not a member?
      1. (a)
        France
      2. (b)
        China
      3. (c)
        India
    • Question of the IMF. Where are the headquarters of the International Monetary Fund (IMF) located?
      1. (a)
        Washington DC
      2. (b)
        London
      3. (c)
        Geneva
  4. 4.
    Ethical norms and laws
    “Please tell me for each of the following actions whether you think it can always be justified, never be justified, or something in between.
    1 = Never justifiable, 2 , 3 , 4 , 5 , 6 , 7 , 8 , 9 , 10 = Always justifiable”
    • Fare evasion acceptance. Avoiding a fare on public transport
    • Theft acceptance. Stealing property
    • Tax evasion acceptance. Cheating on taxes if you have a chance
    • Bribery acceptance. Someone accepting a bribe in the course of their duties
    • Homosexuality acceptance. Homosexuality
    • Divorce acceptance. Divorce
    • Parental abuse acceptance. Parents beating children
    • Terrorism acceptance. Terrorism as a political, ideological or religious mean
  5. 5.
    Out-of-sample questions in Figure 4
    “For each of the following statements, can you tell me how strongly you agree or disagree with each. Do you strongly agree, agree, disagree, or strongly disagree? Choose 1 for Strongly agree, 2 for Agree, 3 for Disagree, and 4 for Strongly disagree.”
    • Parental pride goal. One of my main goals in life has been to make my parents proud
    • Work-motherhood impact. When a mother works for pay, the children suffer
    • Male political leadership. On the whole, men make better political leaders than women do
    • University gender gap. A university education is more important for a boy than for a girl
    • Male business executives. On the whole, men make better business executives than women do
    • Housewife fulfillment. Being a housewife is just as fulfilling as working for pay
LLM prompt: ANES
Step 1: Assigning a persona to the LLM
“It is the year 2024. You are a/an [AGE]-year-old [ETHIC GROUP] [GENDER] living in the United States, who are registered to vote in [STATE], with [EDUCATION LEVEL] education level and you [RELIGIOUS PHRASE]. Your marital status is [MARITAL STATUS], and you [OCCUPATION DESCRIPTION], [HOW OFTEN] paying attention to what’s going on in government and politics. The income of all members of your family during the past 12 months before taxes is [INCOME LEVEL].”
Step 2: Setting the scenario
“Hello!
The presidential election plays a crucial role in determining the direction of the United States for the next four years, influencing not only domestic policies like the economy and healthcare but also having a profound impact on global affairs.
Every vote matters in shaping the future of the nation, as participating in the election is not only a right but also a responsibility.
We encourage everyone to make decisions based on policies and stances rather than emotions or stereotypes.
Now we are conducting a scientific study involving a mock election for the U.S. presidential election.
Assume that you are a voter participating in the 2024 U.S. election. You will be asked to cast your votes. After voting, the results will be shared. Your voting results will only be used for academic analysis, so please don’t worry and feel free to vote.
U.S. Presidential and Vice Presidential Election - Vote for One Party
Step 3: Presenting options and gathering responses
Please select the presidential and vice presidential candidates you support:
  1. 1.
    KAMALA D. HARRIS / TIM WALZ (Democratic)
  2. 2.
    DONALD J. TRUMP / J.D. VANCE (Republican)
Note: Each voter can only select one party ticket. The party label accompanying the candidates indicates that they are the official nominees of the party shown.
Please select your choice: 1 or 2. Respond only with the corresponding number.”
The presidential and vice presidential candidates were replaced with the following if the year was set to 2016 or 2020.
  • 2016 United States presidential election
    1. 1.
      HILLARY R. CLINTON / TIMOTHY M. KAINE (Democratic)
    2. 2.
      DONALD J. TRUMP / MICHAEL R. PENCE (Republican)
  • 2020 United States presidential election
    1. 1.
      JOSEPH R. BIDEN / KAMALA D. HARRIS (Democratic)
    2. 2.
      DONALD J. TRUMP / MICHAEL R. PENCE (Republican)
Appendix A Demographics
For each query to ChatGPT, the bracketed characteristics are replaced with values matching those of an actual respondent from Wave 7 of the WVS. A detailed list of these values is provided as follow:
Demographics Descriptions
AGE age in years
GENDER male or female
STATE/PROVINCE a U.S. state or a Chinese province
EDUCATION LEVEL 0: an early childhood education level,
1: a primary education level,
2: a lower secondary education level,
3: an upper secondary education level,
4: a post-secondary non-tertiary education level,
5: a short-cycle tertiary education level,
6: a bachelor or equivalent education level,
7: a master or equivalent education level,
8: a doctoral or equivalent education level
MARITAL STATUS 1: married,
2: living together as married,
3: divorced,
4: separated,
5: widowed,
6: single
OCCUPATION DESCRIPTION 1: work in a professional and technical field (for example: doctor, teacher, engineer, artist, accountant, nurse),
2: work in higher administrative (for example: banker, executive in big business, high government official, union official),
3: work in clerical (for example: secretary, clerk, office manager, civil servant, bookkeeper),
4: work in sales (for example: sales manager, shop owner, shop assistant, insurance agent, buyer),
5: work in service (for example: restaurant owner, police officer, waitress, barber, caretaker),
6: work as a skilled worker (for example: foreman, motor mechanic, printer, seamstress, tool and die maker, electrician),
7: work as a semi-skilled worker (for example: bricklayer, bus driver, cannery worker, carpenter, sheet metal worker, baker),
8: work as an unskilled worker (for example: laborer, porter, unskilled factory worker, cleaner),
9: work as a farm worker (for example: farm laborer, tractor driver),
10: work as a farm proprietor, farm manager,
11: are retired/pensioned,
12: are a housewife not otherwise employed,
13: are a student,
14: are unemployed
INCOME LEVEL 0 (the lowest income group) to 10 (the highest income group).
For each query to ChatGPT, the bracketed characteristics are replaced with values matching those of an actual respondent from the ANES. A detailed list of these values is provided as follow:
Demographics Descriptions
AGE age in years
GENDER male or female
ETHIC GROUP 1: non-Hispanic white,
2: non-Hispanic black,
3: Hispanic,
4: non-Hispanic Asian or Native Hawaiian/other Pacific Islander,
5: non-Hispanic Native American/Alaska Native or other race,
6: non-Hispanic of multiple races
STATE/PROVINCE a U.S. state
EDUCATION LEVEL 1: a less than high school credential,
2: a high school diploma or equivalent,
3: a some college but no degree,
4: an associate degree in college(occupational/vocational),
5: an associate degree in college(academic),
6: a bachelors degree,
7: a masters degree,
8: a professional school degree / doctoral degree
RELIGIOUS PHRASE 1: belong to the Protestant faith,
2: belong to the Roman Catholic faith,
3: belong to the Orthodox Christian (such as Greek or Russian Orthodox) faith,
4: belong to the Latter-Day Saints(LDS) faith,
5: belong to the Jewish faith,
6: belong to the Muslim faith,
7: belong to the Buddhist faith,
8: belong to the Hindu faith,
9: belong to the Atheist faith,
10: belong to the Agnostic faith,
11: belong to a minority religious group,
12: do not belong to a denomination
MARITAL STATUS 1: married(spouse present),
2: married(spouse absent),
3: widowed,
4: divorced,
5: separated,
6: never married
OCCUPATION DESCRIPTION 1: work in a for-profit company or organization,
2: work in a non-profit organization (including tax-exempt and charitable organizations),
3: work in local government (for example: city or county school district),
4: work in state government (including state colleges/universities),
5: serve on active duty U.S. Armed Forces or Commissioned Corps,
6: work as a federal government civilian employee,
7: work as an owner of non-incorporated business, professional practice, or farm,
8: work as an owner of incorporated business, professional practice, or farm,
9: work without pay in a for-profit family business or farm for 15 hours or more per week
HOW OFTEN 1: always,
2: most of the time,
3: about half the time,
4: some of the time,
5: never
INCOME LEVEL The income level variable has 22 categories, ranging from “under \$9,999” to “\$250,000 or more” with intervals of approximately \$5,000 to \$25,000.
Appendix B Tables
Table 3: State-Level LLM-predicted 2024 Election
State Role-play prompt ($\hat{h}=0.8$) Structural prompt Polling
Democrats Republicans Democrats Republicans Democrats Republicans
(1) (2) (3) (4) (5) (6)
Alabama 34.79% 61.01% 13.22% 86.78%
Alaska 32.28% 60.81% 0.00% 100.00% 43.00% 51.00%
Arizona 42.71% 52.57% 32.60% 67.40% 46.80% 49.00%
Arkansas 32.87% 62.50% 6.06% 93.94% 40.00% 55.00%
California 62.73% 32.42% 97.60% 2.40% 59.00% 34.30%
Colorado 49.79% 43.25% 74.06% 25.94%
Connecticut 55.48% 41.95% 97.41% 2.59%
Delaware 53.01% 40.48% 96.58% 3.42%
Florida 45.08% 51.57% 20.78% 79.22% 44.20% 51.40%
Georgia 47.58% 49.26% 50.52% 49.48% 46.80% 48.80%
Hawaii 63.13% 31.65% 100.00% 0.00%
Idaho 27.45% 63.80% 0.00% 100.00%
Illinois 54.90% 40.95% 91.37% 8.63%
Indiana 35.66% 59.70% 3.51% 96.49%
Iowa 38.41% 56.92% 3.15% 96.85% 44.00% 49.00%
Kansas 36.40% 58.28% 2.53% 97.47% 43.00% 48.00%
Kentucky 33.15% 62.31% 1.76% 98.24%
Louisiana 38.07% 58.98% 18.25% 81.75%
Maine 46.96% 46.57% 68.25% 31.75% 48.00% 41.00%
Maryland 63.45% 32.39% 97.68% 2.32% 61.30% 33.00%
Massachusetts 62.74% 32.71% 99.75% 0.25% 60.50% 32.00%
Michigan 45.91% 49.27% 54.99% 45.01% 48.90% 47.10%
Minnesota 47.62% 46.38% 75.73% 24.27% 50.00% 43.70%
Mississippi 42.05% 54.96% 35.03% 64.97%
Missouri 35.07% 60.21% 5.63% 94.37% 42.00% 53.50%
Montana 31.53% 63.00% 0.00% 100.00% 39.50% 57.50%
Nebraska 29.66% 65.69% 0.00% 100.00% 48.40% 47.20%
Nevada 49.93% 45.60% 78.78% 21.22% 47.70% 48.10%
New Hampshire 44.18% 49.96% 58.91% 41.09% 50.30% 44.00%
New Jersey 55.24% 41.88% 93.65% 6.35% 52.00% 40.00%
New Mexico 51.47% 41.37% 81.11% 18.89% 49.70% 42.70%
New York 58.67% 37.26% 97.32% 2.68% 57.50% 39.00%
North Carolina 45.05% 50.55% 27.76% 72.24% 47.20% 48.60%
North Dakota 23.88% 68.07% 0.37% 99.63%
Ohio 39.23% 56.24% 11.30% 88.70% 45.00% 51.40%
Oklahoma 29.07% 65.07% 0.61% 99.39% 40.00% 56.00%
Oregon 52.93% 39.80% 97.47% 2.53% 53.00% 41.00%
Pennsylvania 45.81% 50.18% 44.16% 55.84% 48.00% 48.60%
Rhode Island 49.86% 47.23% 98.10% 1.90% 57.00% 40.50%
South Carolina 41.72% 53.92% 18.93% 81.07% 42.00% 53.50%
South Dakota 30.46% 63.21% 0.00% 100.00% 34.00% 60.50%
Tennessee 34.34% 61.88% 6.48% 93.52% 35.00% 56.00%
Texas 43.05% 53.12% 12.17% 87.83% 45.20% 51.40%
Utah 31.16% 56.06% 1.57% 98.43% 38.00% 54.00%
Vermont 63.41% 32.12% 100.00% 0.00% 70.00% 29.00%
Virginia 49.92% 45.91% 53.66% 46.34% 50.00% 41.30%
Washington 53.33% 38.87% 96.51% 3.49% 56.30% 35.70%
Washington DC 87.52% 9.39% 100.00% 0.00%
West Virginia 24.60% 70.43% 0.00% 100.00% 34.00% 61.00%
Wisconsin 44.29% 50.97% 41.89% 58.11% 48.10% 48.30%
Wyoming 23.80% 72.08% 0.00% 100.00%